# 1]FAIR at Meta

### v1

@@ 4 Training and searching partitions
In this section, we present more details how our Saap method assigns keys and queries to buckets. Note already, a key will be assigned to a single partition bucket (unlike LSH). Queries are assigned to the $\ell$ most promising buckets, in a multi-probe fashion.
4.1 Partitioning functions
The assignment function aims at assigning key vectors to a relevant bucket and finding the $-\ell$ best buckets for the queries. A common solution is to partition the data with LSH or clustering. Clustering is more effective than LSH variants (see Appendix 7), that are not data-adaptive. Thus, we use the standard k-means partition (Jegou et al., 2010) by clustering $K\in\mathbb{R}^{n_{\mathrm{k}}\times d}$ keys into $C$ centroids. K-means has the advantage to produce relatively balanced clusters.
Assigning with RoPE.
The first issue we run into with this approach is the temporal OOD behavior, meaning that long-range lookups that are relatively independent of position are hard to match. A large part of this is due to the RoPE transformation of vectors. Therefore, we apply the k-means clustering on vectors before transforming them with RoPE (Section 3.4). The keys and queries from which the attention is computed are still transformed with RoPE.
Asymmetric assignment.
The second adaptation we do is to address the key-query OOD setting (see Section 3). For this we separate the assignment function for keys and queries: for queries we train a classifier $f_{\mathrm{q}}$ that outputs a probability distribution over $[0,1]^{C}$. At search time, for query $q$, we compute the attention only for the keys in the clusters given by $\mathrm{argmax}_{\ell}f_{\mathrm{q}}(q)$.
4.2 Query assignment function
Here we assume that k-means is already trained. We use a training set of queries $Q\in\mathbb{R}^{n_{\mathrm{k}}\times d}$ and keys $K\in\mathbb{R}^{n_{\mathrm{k}}\times d}$. The hard assignment is represented as a binary matrix $H_{\mathrm{k}}=\{0,1\}^{n_{\mathrm{k}}\times C}$, where entry $(i,j)$ is set to 1 if key $i$ was assigned to bucket $j$.
Entry $(i,j)$ of the matrix $A\times H_{\mathrm{k}}$ represents the fraction of attention weight for query $i$ contained in bucket $j$. Therefore, the output distribution of $f_{\mathrm{q}}$ should be as similar as possible to row $i$ of $A\times H_{\mathrm{k}}$. This yields the following loss:
@@ 4.2 Query assignment function
where $\mathrm{KLDiv}$ is the Kullback-Leibler divergence between distributions, $\mathrm{KLDiv}$ and $f_{\mathrm{q}}$ are applied row-wise on their matrix arguments. We optimize this function using AdamW Loshchilov et al. (2017). Each training batch comes from a separate prompt. We sample $n_{\mathrm{q}}$ queries $Q$ and $n_{\mathrm{k}}$ keys $K$ uniformly from the prompt. Since short-term queries are taken into account by the “1+2047” dense part, we force the training to focus on long range queries by sampling only (key, query) pairs that are more than 2047 steps apart.
4.3 Beyond top-k
The standard IVF approach retrieves only the top-$k$ keys that have the largest dot product with the query. However, collecting the top-$k$ results in vector search is surprisingly slow (Johnson et al., 2019), especially on GPUs. Besides, retrieving the top $k$ does require to compute the dot products w.r.t. all keys within a matching bucket anyways. Therefore, we aggregate all the corresponding values into the attention result on-the-fly.
4.4 Implementation
Saap focuses on the decoding phase of the generation because this is when the computation is most egregiously slowed by memory access times. An efficient implementation requires batched queries and balanced clusters.
Batched queries.
Llama 3 8B models use group query attention (Ainslie et al., 2023), where four heads from one layer share the same KV cache. At search time, for each token, we need to perform 4 searches for queries $\{q_{1},q_{2},q_{3},q_{4}\}$. For efficiency, we perform the 4 searches jointly by combining the outputs of the query assignment model as follows:
$\mathrm{argmax}_{\ell}\sum_{i=1}^{4}f_{\mathrm{q}}(q_{i}).$ (2)
This approximation converts a matrix-vector into a matrix-matrix multiplication (albeit with a small batch size of 4). In Section 5.5 we measure the impact of this approximation.
On GPU.
The Saap kernel follows an implementation similar to the Triton-based Flash-Decoding (Dao et al., 2023) included in xFormers (Lefaudeux et al., 2022). Contrary to Flash-Decoding, which splits the keys and values uniformly across different CUDA blocks, in our implementation each CUDA block computes the partial attention for a single cluster assignment, with the dense local attention (Section 3.5) happening on its own CUDA block. The custom kernel is necessary in order to limit expensive synchronizations between CPU and GPUs due to the variable sizes of the visited clusters. It could benefit from further optimizations as devised by Shah et al. (2024).
The issue with this approach is that the runtime for one layer depends on the largest cluster that is accessed by any of the heads in the layer (i.e. the maximum over $8\times\ell$ clusters). Therefore, we rely on the key partitioning to produce balanced clusters. The approach chosen by KDEFormer and Reformer is to artificially force the clusters to be balanced, but with a significant impact on accuracy (see Appendix 7).
@@ 5 Experiments
In this section, we evaluate Saap on two natural language processing (NLP) benchmarks and report end-to-end LLM performance metrics. We provide a comparison with the more important baselines. Note, Appendix 7 provides an additional comparison with LSH partitioning.
Figure 4: Scores of Needle-in-haystack (left) and the En.QA task from InfiniteBench (right) as a function of the budget (fraction of KV-cache visited), on a Llama 3 8B with scaled rope to 500k. We compare several ways of assigning keys and queries to buckets. We also show the score with no sparse attention (StreamingLLM) and the full attention score.
@@ 5.1 Models, measures, benchmarks and baselines
We perform experiments using a Llama 3.1 8B, a LLM with 32 layers, 32 query heads and 8 key heads. This model has been trained for context sizes up to 128k tokens. We extend the context size to 500k tokens using rope scaling (Peng et al., 2023). We do not fine-tune the model weights.
@@
Saap first partitions the keys with a pre-trained k-means and trains a Multilayer Perceptron (MLP) in order to assign queries to buckets (offline), denoted by K-means+Q-model. The MLP has two layers with an intermediate dimension of 1024, batch normalization and a ReLU nonlinearity. We use the Saap assignment models on PG19 books (Rae et al., 2019) for the MLP training. These contain natural-language text, but may be out-of-distribution for math or coding tasks. We sample key and query vectors across the prompt for training, and keep only long-range queries: we discard all queries whose distance to the top-1 matching key is less than 1024 tokens. We use frozen key and query embeddings for training. The training takes around 25 minutes per head on one GPU. This is not a limitation since the training is meant to be re-used across contexts and prompts.
Performance metrics: selectivity and generation timings.
@@
The main goal of Saap is to speed up the attention computation and, consequently, decoding. Therefore, we measure and report resource consumption as a trade-off between accuracy and computing speed. We measure the computing speed in two ways: the first one is the selectivity (Chen et al., 2024b; Paulevé et al., 2010), or the percentage of keys that are involved in the computation, averaged over all heads; the second one is a direct timing of the end-to-end generation. We used H100 GPUs to measure the timings.
Benchmarks.
@@
We evaluate the end-to-end accuracy on long-context tasks that require different types of lookups (Li et al., 2024a), more precisely:
@@
  • Needle-in-a-Haystack (Kamradt, 2023) is a benchmark that consists on a “hard” text matching task where a secret, short text (the needle) is inserted at an arbitrary position in an arbitrarily long context made of book texts (which are in-domain for the training data). The prompt then attempts to retrieve the text. We average the Needle-in-a-Haystack scores over prompt lengths between 10k and 128k, reporting separately for lengths between 128k and 500k, and for various insertion points for the needle.
@@
  • InfiniteBench benchmark (Zhang et al., 2024) is a series of tasks intended to probe LLM reasoning capabilities beyond a 100k context size. The benchmarks relate to question answering, coding and math questions as well as retrieval tasks for numbers, passkeys and KV.
Baselines.
We compare Saap to the following methods:
@@
  • Full attention is the default brute-force attention computation. We use FlashAttention-v2 (Dao, 2024), which is a strong open-source reference implementation for long-context attention.
@@
  • K-means refers to a partition-based approximate attention using k-means partitioning. We report results with k-means directly applied to the full keys and queries, as well as applied to keys and queries without the rope transformation. This is similar to Faiss’ IVFFlat index (Douze et al., 2024), that has been used as a baseline by Liu et al. (2024a), but not restricted to the top-k nearest keys (see Section 4.4).
@@
  • StreamingLLM Xiao et al. (2024c) this corresponds to setting $\ell=0$ in IVFFlat, using only the dense part of the attention computation.
@@ 5.2 Comparison with baselines
Table 2: Scores with a Llama 3.1-8B model using rope extrapolation to increase the context length of 500k and different partitioning schemes. We report in bold the best results among approximate attention schemes.
Methods NIH InfiniteBench (Zhang et al., 2024)
@@
K-means (roped inputs, $\ell=32$) 95.48 69.33 98.98 100.00 4.40 26.00 31.47 25.42 48.03
K-means (no rope, $\ell=32$) 100.00 98.67 99.66 100.00 4.40 34.29 31.47 29.58 50.66
@@
Saap, K-means+Q ($\ell=32$) 100.00 100.00 100.00 100.00 7.40 32.57 30.20 29.68 50.21
Saap, K-means+Q ($\ell=32$, selectivity) 4.0 5.3 15.0 13.8 21.9 47.8 4.8 3.7 3.7
Table 3: Comparison of Saap with a few other methods from the state of the art. The base model is a Llama 3 8B finetuned by Gradient AI for context length of 262k. Rows marked with * are reported by Liu et al. (2024a).
InfiniteBench (Zhang et al., 2024)
Methods Retr.N Retr.P Retr.KV Math.F Code.D En.QA En.MC
@@
Saap, K-means+Q ($\ell=$ 32) 100.0 100.0 5.5 38.5 25.5 11.6 56.0
Saap, K-means+Q ($\ell=$ 32, selectivity) 25.8 25.9 19.1 51.0 4.8 4.3 4.3
Figure 4 shows the trade-off between budget and score (we report detailed numbers in Table 2). The operating point is controlled by setting the query-time $\ell\in\{0,4,8,16,32,64,128\}$. We also show 1) the $\ell=0$ setting (StreamingLLM  Xiao et al. (2024c)), that corresponds to the bottom line; and 2) the full attention, where the keys are compared exhaustively, the topline.
Needle-in-a-Haystack scores reach 100% accuracy from $\ell$$=$$8$ for context sizes up to 128k, but requires $\ell$$=$$32$ beyond that. On InfiniteBench, the behavior depends on the tasks. In Figure 4, En.QA exhibits is a clear positive trend, the topline accuracy is reached with around 5% selectivity.
For Code.D and En.MC the results for $\ell=0$ and full attention are close, so the scores are noisy. For Math.F, the selectivity increases quickly with $\ell$, which is indicative of an OOD behavior between training and testing prompts. See also  Figure 9 in  section 9. This is because for this task, the context consists of a long lists of numbers, which is very different from the books on which the assignment models are trained. This causes a large imbalance factor.
Interestingly, the Saap score is higher than the full attention score. This may be because the sparse attention focuses only on a relevant subset of the prompt instead of averaging information from the entire sequence.
@@ 5.3 Timings
End-to-end timings do not only depend on the budget, but also on how balanced the clusters are and whether the heads all perform the same number of FLOPs. This is due to the parallelization over GPUs and over computation cores that we use (Section 4.4).
We measure the attention computation speed for one head on a 171k context (with $\ell=32$, yielding 100% needle-in-a-haystack accuracy). We compare against a strong attention implementation, FlashAttention-v2 Dao et al. (2023).

### v2

@@ 4 Training and searching partitions
In this section, we present more details on how our Saap method assigns keys and queries to buckets. Note that a key will be assigned to a single partition bucket (unlike LSH). On the other hand, queries are assigned to the $\ell$ most promising buckets, in a multi-probe fashion. The assignment function aims at assigning key vectors to a relevant bucket and finding the $-\ell$ best buckets for the queries. A common solution is to partition the data with LSH or clustering. Clustering is more effective than LSH variants (see Appendix B), which are not data-adaptive. Thus, we use standard k-means (Jegou et al., 2010) for clustering $K\in\mathbb{R}^{n_{\mathrm{k}}\times d}$ keys into $C$ centroids. K-means has the advantage that it produces relatively balanced clusters.
Asymmetric assignment.   The first adaptation we perform is to address the key-query OOD setting (see Section 3). For this, we separate the assignment function for keys and queries: for queries we train a classifier $f_{\mathrm{q}}$ that outputs a probability distribution over $[0,1]^{C}$. At search time, for query $q$, we compute the attention only for the keys in the clusters given by $\mathrm{argmax}_{\ell}f_{\mathrm{q}}(q)$.
Assigning with(out) RoPE.    The second issue with KV cache indexing is the temporal OOD behavior, meaning that long-range lookups that are relatively independent of position are hard to match. A large part of this is due to the RoPE transformation of vectors. Therefore, we apply the k-means clustering on vectors before transforming them with RoPE. The keys and queries from which the attention is computed are still transformed with RoPE.
Query assignment function   Here we assume that k-means is already trained. We use a training set of queries $Q\in\mathbb{R}^{n_{\mathrm{k}}\times d}$ and keys $K\in\mathbb{R}^{n_{\mathrm{k}}\times d}$. The hard assignment is represented as a binary matrix $H_{\mathrm{k}}=\{0,1\}^{n_{\mathrm{k}}\times C}$, where entry $(i,j)$ is set to 1 if key $i$ was assigned to bucket $j$. Then, the entry $(i,j)$ of the matrix $A\times H_{\mathrm{k}}$ represents the fraction of attention weight for query $i$ contained in bucket $j$. Therefore, the output distribution of $f_{\mathrm{q}}$ should be as similar as possible to row $i$ of $A\times H_{\mathrm{k}}$. This yields the following loss:
@@
where $\mathrm{KLDiv}$ is the Kullback-Leibler divergence between distributions, and $\mathrm{KLDiv}$ and $f_{\mathrm{q}}$ are applied row-wise on their matrix arguments. Each training batch comes from a separate prompt. We sample $n_{\mathrm{q}}$ queries $Q$ and $n_{\mathrm{k}}$ keys $K$ uniformly from the prompt. Since short-term queries are taken into account by the “1+2047” dense part, we force the training to focus on long range queries by sampling only (key, query) pairs that are more than 2047 steps apart. See Appendix A for details about the training parameters.
Beyond top-k.    The IVF approach retrieves only the top-$k$ keys that have the largest dot product with the query. However, collecting the top-$k$ results in vector search is slow (Johnson et al., 2019) on GPUs. Besides, retrieving the top $k$ does require computing the dot products w.r.t. all keys within a matching bucket anyways. Therefore, we aggregate all the corresponding values into the attention result on-the-fly.
5 Implementation
Batched queries.   Llama 3 8B models use group query attention (Ainslie et al., 2023), where four heads from one layer share the same KV cache. At search time, for each token, we need to perform 4 searches for queries $\{q_{1},q_{2},q_{3},q_{4}\}$. For efficiency, we perform the 4 searches jointly by combining the outputs of the query assignment model as $\sum_{i=1}^{4}f_{\mathrm{q}}(q_{i})\in\mathbb{R}^{C}$, and pick the $-\ell$ values from this vector. This approximation converts a matrix-vector product into a matrix-matrix multiplication (albeit with a small batch size of 4). In Section 6.5 we measure the impact of this approximation.
On GPU.   The Saap kernel follows an implementation similar to the Triton-based Flash-Decoding (Dao et al., 2023) included in xFormers (Lefaudeux et al., 2022). In contrast to Flash-Decoding, which splits the keys and values uniformly across different CUDA blocks, in our implementation, each CUDA block computes the partial attention for a single cluster assignment, with the dense local attention (Section 2) occurring in its own CUDA block. The custom kernel is necessary in order to limit expensive synchronizations between CPU and GPUs due to the variable sizes of the visited clusters. It could be further optimized in a way to similar to Shah et al. (2024).
The issue with this approach is that the runtime for one layer depends on the largest cluster that is accessed by any of the heads in the layer (i.e. the maximum over $8\times\ell$ clusters). Therefore, we rely on the key partitioning to produce balanced clusters. Note that the KDEFormer and Reformer enforce the clusters to be artificially balanced, this has a significant impact on accuracy (see Appendix B).
@@ 6 Experiments
In this section, we evaluate Saap on two natural language processing (NLP) benchmarks and report end-to-end LLM performance metrics. We provide a comparison with the most important baselines. Appendix B provides an additional comparison with LSH partitioning, while Appendices C and D report additional results on the Ruler (Hsieh et al., 2024) and InfiniteBench (Zhang et al., 2024b) benchmarks, respectively.
Figure 4: Scores of Needle-in-haystack (left) and the En.QA task from InfiniteBench (right) as a function of the selectivity (fraction of KV-cache visited), on a Llama 3 8B with scaled rope to 500k. We compare several ways of assigning keys and queries to buckets. We also show the score with no sparse attention (StreamingLLM) and the full attention score.
@@ 6.1 Models, measures, benchmarks and baselines
We perform experiments using a Llama 3.1 8B, a LLM with 32 layers, 32 query heads and 8 key heads. This model is trained for context sizes up to 128k tokens. We extend this size to 500k tokens using rope scaling (Peng et al., 2023). Our method does not require to fine-tune the model weights.
@@
Saap first partitions the keys with a pre-trained k-means and trains a Multilayer Perceptron (MLP) in order to assign queries to buckets (offline), denoted by K-means+Q-model. The MLP has two layers with an intermediate dimension of 1024, batch normalization and a ReLU nonlinearity. We use the Saap assignment models on the PG19 books dataset (Rae et al., 2019) for training the MLP. This data contains natural-language text, but may be out-of-distribution for math or coding tasks. We sample key and query vectors across the prompt for training, and keep only long-range queries: we discard all queries whose distance to the top-1 matching key is less than 1024 tokens. We use frozen key and query embeddings for training. See Appendix A for more training details.
@@
Performance metrics: selectivity and generation timings.    The main goal of Saap is to speed up the attention computation and, consequently, decoding. Therefore, we measure and report resource consumption as a trade-off between accuracy and computing speed. We measure the computing speed in two ways: the first is selectivity (Chen et al., 2024b; Paulevé et al., 2010), or the rate of keys that are involved in the computation, averaged over all heads; the second is a direct timing of the end-to-end generation.
@@
Benchmarks.   We evaluate the end-to-end accuracy on long-context tasks that require different types of lookups (Li et al., 2024a), more precisely:
@@
  • Needle-in-a-Haystack (Kamradt, 2023) consists of a “hard” text matching task where a secret, short text (the needle) is inserted at an arbitrary position in an arbitrarily long context made of book texts (in-domain for the training data). The prompt attempts to retrieve the text. We average the scores over prompt lengths between 10k and 128k, reporting separately for lengths between 128k and 500k, and for various insertion points for the needle.
@@
  • InfiniteBench benchmark (Zhang et al., 2024b) is a series of tasks designed to probe LLM reasoning capabilities beyond a context size of 100k. The tasks relate to question answering, coding and math questions, as well as retrieval tasks for numbers, passkeys and KV.
Baselines.   We compare Saap to the following methods:
@@
  • Full attention is the default brute-force attention computation. We use FlashAttention-v2 (Dao, 2024), a strong open-source reference implementation for long-context attention.
@@
  • K-means refers to a partition-based approximate attention using k-means partitioning. We report results with k-means directly applied to the full keys and queries, as well as applying it to keys and queries without the rope transformation. This is similar to Faiss’ IVFFlat index (Douze et al., 2024), which has been used as a baseline by Liu et al. (2024a), but in our case it is not restricted to the top-k nearest keys (see Section 4: beyond top-k).
@@
  • StreamingLLM (Xiao et al., 2024b) corresponds to setting $\ell=0$ in IVFFlat, using only the dense part of the attention computation.
@@ 6.2 Comparison with baselines
Table 2: Scores with a Llama 3.1-8B model using rope extrapolation to increase the context length of 500k and different partitioning schemes. We report in bold the best results among approximate attention schemes. All methods involving k-means assignment use $\ell$ = $32$ for multi-probing.
Method NIH InfiniteBench (Zhang et al., 2024b)
@@
K-means (roped inputs) 95.48 69.33 98.98 100.00 4.40 26.00 31.47 25.42 48.03
K-means (no rope) 100.00 98.67 99.66 100.00 4.40 34.29 31.47 29.58 50.66
@@
Saap, K-means + Q 100.00 100.00 100.00 100.00 7.40 32.57 30.20 29.68 50.21
selectivity $\rightarrow$ 4.0 5.3 15.0 13.8 21.9 47.8 4.8 3.7 3.7
Table 3: Comparison of Saap ($\ell$ = $32$) with state of the art methods. The base model is a Llama 3 8B finetuned by Gradient AI for context length of 262k. *: rows reported from Liu et al. (2024a).
InfiniteBench (Zhang et al., 2024b)
    Method Retr.N Retr.P Retr.KV Math.F Code.D En.QA En.MC
@@
    Saap, K-means + Q 100.0 100.0 5.5 38.5 25.5 11.6 56.0
    selectivity $\rightarrow$ 25.8 25.9 19.1 51.0 4.8 4.3 4.3
Figure 4 shows the trade-off between budget and score (we report detailed numbers in Table 2). The operating point is controlled by setting the query-time $\ell\in\{0,4,8,16,32,64,128\}$. We also show 1) the $\ell=0$ setting (StreamingLLM  (Xiao et al., 2024b)), that corresponds to the bottom line; and 2) the full attention, where the keys are compared exhaustively, as the topline.
Needle-in-a-Haystack scores reach 100% accuracy from $\ell$$=$$8$ for context sizes up to 128k, but requires $\ell$$=$$32$ beyond that. On InfiniteBench, the behavior depends on the tasks. In Figure 4, En.QA exhibits a clear positive trend, and the topline accuracy is reached with around 5% selectivity.
For Code.D and En.MC the results for $\ell=0$ and full attention are close, so the scores are noisy. For Math.F, the selectivity increases quickly with $\ell$, which is indicative of an OOD behavior between training and testing prompts. See also  Figure 6 in  Appendix D. This is because for this task, the context consists of long lists of numbers, which is very different from the books on which the assignment models are trained. This causes a large imbalance factor. Interestingly, the Saap score is higher than the full attention score. This may be because the sparse attention focuses only on a relevant subset of the prompt instead of averaging information from the entire sequence.
@@ 6.3 Timings
Generation. We use 8 H100 GPUs 80 GB (1 KV-cache per GPU) and measure the attention computation speed for one head on a 171k-length context ($\ell$ = $32$, yielding 100% needle-in-a-haystack accuracy). We compare against the FlashAttention-v2 implementation (Dao et al., 2023).